Q1) Solve the following DE: (Using Laplace transform is recommended) y + 5y – 6y = f(t), y(0) = 0, y(0) = 0, where 0 <t< 2
Answer #1

y + 54-6y=flt) Y(0)= 0, y(o)=0 oster f(t) = si let & 7,2 Note that Lsy}= 3² y(s) - syco)-y (o) and L{y} = 8.4 (s) - glo)1 t 7,2 st1 1 1 + B s-1 + S+6 and – 6A = 1 Y(3) (8² +58-6) = ost 2 for 0<432!- Y(3) 1 s(3²+5s-6) 8(3-1) (3+6) Write A S(S-1)and to get! Put A = -16 in B+c= 172 3 and C-6B = 15 tolo or 6B-C-5 6 6 Add 9 and 14 to get :- 7B = 6 6 1 = B = 3/4 Thus, (= 2Y(8) = 1 (8 + 1) (8² +58-) 1 (8+1) (3-1)(3+6) 1 А St1 + B s-1 + (8+1) (8-1)(+6) st с sto (8+1) (8-1) (3+6) (1-1)(3+6) A + (+130 B+ 420-2B 5 70 14 A= -73 = -1/2 and c= 2B 3 35 Thus, (-430), (434) + (+135) 1 (SH) (87)(3+6) + St1 s-1 s+6 So, Y(s) = (1/1